# 4D Radar-Inertial Odometry based on Gaussian Modeling and Multi-Hypothesis Scan Matching

### v4

@@ Abstract
4D millimeter-wave (mmWave) radars are sensors that provide robustness against adverse weather conditions (rain, snow, fog, etc.), and as such they are increasingly used for odometry and SLAM (Simultaneous Location and Mapping). However, the noisy and sparse nature of the returned scan data proves to be a challenging obstacle for existing registration algorithms, especially those originally intended for more accurate sensors such as LiDAR. Following the success of 3D Gaussian Splatting for vision, in this paper we propose a summarized representation for radar scenes based on global simultaneous optimization of 3D Gaussians as opposed to voxel-based approaches, and leveraging its inherent Probability Density Function (PDF) for registration. Moreover, we propose tackling the problem of radar noise entirely within the scan matching process by optimizing multiple registration hypotheses for better protection against local optima of the PDF. Finally, following existing practice we implement an Extended Kalman Filter-based Radar-Inertial Odometry pipeline in order to evaluate the effectiveness of our system. Experiments using publicly available 4D radar datasets show that our Gaussian approach is comparable to existing registration algorithms, outperforming them in several sequences.
@@ I Introduction
Figure 1: Overview of the Radar-Inertial Odometry pipeline including our proposed Gaussian modeling and scan matching processes. Bold font indicates input/output, the highlighted area indicates our contribution
We introduce a new method for modeling and registering 4D radar point clouds using 3D Gaussians. Following the success of 3D Gaussian Splatting (3DGS) for vision [6], we propose the use of a collection of 3D Gaussians for modeling the geometry of a scene. As in [6] (and in contrast to NDT [1]), we simultaneously optimize all parameters of the model, including the position, scaling, rotation and correspondences between points and Gaussians. This approach is entirely featureless, and does not require any priors other than an initial set of centroids. Moreover, we propose a registration step that considers multiple simultaneous hypotheses in order to add further robustness against radar noise. Finally, we integrate our algorithms into a Radar-Inertial Odometry (RIO) pipeline using an Extended Kalman Filter (EKF) (see Fig. 1), which we then use to evaluate our Gaussian methods against several established registration algorithms on multiple datasets and radar types.
This paper is organized as follows: Section II reviews existing related literature, Sections III and IV describe the proposed Gaussian modeling and scan matching components respectively, Section V describes the EKF-based RIO pipeline used in the experiments; Section VI shows experimental results, and Section VII concludes the paper proposing future work.
@@ III-B Initialization
First, it is necessary to determine $N$, that is, the number of 3D Gaussians in the model. We propose choosing $N$ proportionally based on the number of points, so that richer scans can be represented with more Gaussians. The proportion should also consider the sparsity of the point clouds generated by the radar used. The model $\mathcal{G}$ is then initialized with all $\mathbf{s}_{j}$ set to 0 (indicating unit scale), and all $\mathbf{q}_{j}$ set to the unit quaternion. As for $\boldsymbol{\mu}_{j}$, we perform Bisecting K-Means clustering to find initial estimates for the centers of the Gaussians.
@@ III-C Optimization
IV Gaussian multi-hypothesis scan matching
Given a Gaussian model $\mathcal{G}$ and a point cloud $\mathcal{P}$, we define a registration process that estimates the pose of the robot $\xi$ with respect to $\mathcal{G}$. The algorithm is based on the simultaneous optimization of multiple hypotheses centered around the current predicted position of the robot, in order to increase the robustness against local optima. At the end of the optimization, the hypotheses are judged against a scoring function, and the best hypothesis is selected as the output of the scan matching algorithm while the other hypotheses are discarded. Note that this is unlike a classical particle filter, which does persist the hypotheses throughout time.
@@ IV-A Definitions
We define the robot pose $\xi$ to be a member of $SE(3)$; that is, it includes a translational component $\mathbf{t}\in\mathbb{R}^{3}$ and rotational component $\mathbf{q}\in SO(3)$, with respect to some frame of reference. We consider that the uncertainty about the true robot pose follows some probability distribution: $\xi\thicksim\mathrm{P}$. Given a suitable model of $\mathrm{P}$, we can randomly sample a swarm of $K$ pose particles $\hat{\xi}^{k}$, each representing a different hypothesis, and use an optimization algorithm that evaluates each particle against $\mathcal{G}$ to further refine these estimates until one of the particles reaches a suitable optimum. We believe exploring the solution space using a particle swarm increases robustness in challenging situations, especially when dealing with large rotations or translations.
@@ IV-B Optimization
  • We perform the usual Gauss-Newton update step for each particle $\hat{\xi}^{k}$, using the sums of squared Mahalanobis distances as the function to optimize, as in GICP. We also exclude points whose $d^{k}_{i}$ is greater than a given distance threshold $d_{\textrm{max}}$.
@@
$\mathcal{L}_{k}=\frac{1}{M}\sum_{i}^{M}\min(d^{k}_{i},d_{\textrm{max}}),$ (2)
@@
IV-C Keyframing
A common technique used in odometry systems is keyframing. Certain poses of the robot are selected to become local reference points (keyframes) that are typically used for registration, pose graph optimization or (in the case of SLAM) loop closure purposes. In our case, we only use keyframing to select individual radar point clouds to undergo Gaussian modeling, and serve as reference for Gaussian scan matching of subsequent radar point clouds. The system only retains the pose and Gaussian model of the most recently generated keyframe. This approach, compared to trivial successive applications of the scan matching algorithm, has the benefit of reducing the accumulated error, though it also means the criteria used to select keyframe candidates needs to be carefully designed to generate an appropriate keyframe pacing. We use a simple set of criteria to decide when a new pose $\xi^{j}$ qualifies to become a new keyframe. Given the most recent keyframe $\xi^{i}$, we can derive the $SE(3)$ transformation relative to the keyframe $T=\xi^{j}\ominus\xi^{i}=\{\mathbf{t},\mathbf{q}\}$. The criteria are then simply that any of the following must hold:
  • For the translational component: $\mathbf{t}||_{2}\geq{t}_{\max}$, where ${t}_{\max}$ is the maximum translation threshold.
  • For the rotational component: $2\cos^{-1}|q_{w}|\geq\alpha_{\max}$, where $\alpha_{\max}$ is the maximum rotation threshold.
As a fallback, if no scans have returned a successful match during the past given time interval $t$, a keyframe is automatically generated.
V Inertial odometry with EKF
In order to evaluate the suitability and effectiveness of the Gaussian algorithms for Radar-Inertial Odometry, it is necessary to define a pipeline for use in the experiments. Following well established practice [28, 29, 12], we use an Extended Kalman Filter in our odometry solution to tightly integrate all incoming sensor information: inertial data, radar egovelocity observations, and scan matching results. In order to simplify the formulation, we only use Error State filtering for rotational variables, as opposed to the entire state. We employ the right-handed North-West-Up (NWU) convention typical of robotics, and consider three reference frames: world ($w$), body/IMU ($b$) and radar ($r$). The rotation matrix $\mathbf{C}^{g}_{f}$ transforms a vector from frame $f$ to frame $g$, likewise $\mathbf{C}^{f}_{g}=(\mathbf{C}^{g}_{f})^{T}$. We decompose each $\mathbf{C}$ into an error part and a nominal quaternion part as follows: $\mathbf{C}=\exp\left([\delta\boldsymbol{\theta}]_{\times}\right)\mathbf{R}\{\mathbf{q}\}$, approximating the error part as $\mathbf{I}+[\delta\boldsymbol{\theta}]_{\times}$ for the purposes of linearization.
The state ($\mathbf{x}$), control ($\mathbf{u}$) and noise ($\mathbf{w}$) vectors of the filter are the following:
$\mathbf{x}=\begin{bmatrix}\mathbf{p}\\ \mathbf{v}\\ \mathbf{t}_{r}^{b}\\ \mathbf{b}_{a}\\ \mathbf{b}_{\omega}\\ \delta\boldsymbol{\theta}_{b}^{w}\\ \delta\boldsymbol{\theta}_{r}^{b}\end{bmatrix};\;\mathbf{u}=\begin{bmatrix}\mathbf{a}\\ \boldsymbol{\omega}\end{bmatrix};\;\mathbf{w}=\begin{bmatrix}\mathbf{w}_{v}\\ \mathbf{w}_{\theta}\\ \mathbf{w}_{a}\\ \mathbf{w}_{\omega}\\ \mathbf{w}_{b_{a}}\\ \mathbf{w}_{b_{\omega}}\end{bmatrix}$, (3)
where $\mathbf{p}$ is the position, $\mathbf{v}$ is the velocity, $\mathbf{t}_{r}^{b}$ is the translation between radar and IMU, $\mathbf{b}_{a}$ and $\mathbf{b}_{\omega}$ are the accelerometer/gyroscope biases, each $\delta\boldsymbol{\theta}$ is the error part of its corresponding $\mathbf{C}$ frame expressed as a tangential rotation $\in\mathfrak{so}(3)$, and $\mathbf{a}$ and $\boldsymbol{\omega}$ are the IMU accelerometer/gyroscope readings. The position and velocity of the robot are expressed in the world frame, while the nominal rotations $\mathbf{q}_{b}^{w}$ and $\mathbf{q}_{r}^{b}$ are stored outside the filter state as quaternions. Velocity and attitude process noise are modeled by $\mathbf{w}_{v}$ and $\mathbf{w}_{\theta}$ respectively. We follow Kalibr’s IMU noise model [30] [^1] https://github.com/ethz-asl/kalibr/wiki/IMU-Noise-Model, and consider both additive ($\mathbf{w}_{a}$, $\mathbf{w}_{\omega}$) and random walk noise ($\mathbf{w}_{b_{a}}$, $\mathbf{w}_{b_{\omega}}$) in accelerometer/gyroscope readings.
There also exists a vector $\mathbf{g}=-[0\;0\;9.80511]^{T}$ modeling the Earth’s gravity treated as a constant, as we do not consider long enough travel distances for changes in the plane tangent to the Earth’s surface to matter.
V-A EKF initialization
First, we initialize all components of the state vector $\mathbf{x}$ to 0, and the rotational variables to the unit quaternion. The initial radar-to-IMU frame ($\mathbf{t}_{r}^{b}$, $\mathbf{q}_{r}^{b}$) can be supplied externally. The robot is assumed to remain still during the first few seconds, which allows for averaging initial accelerometer/gyroscope readings and extracting an estimate of the body-frame gravity vector $\tilde{\mathbf{g}}_{b}$ (after flipping sign) and the gyroscope bias $\mathbf{b}_{\omega}$ respectively. An initial value for the roll/pitch components of $\mathbf{q}_{b}^{w}$ can then be derived from $\tilde{\mathbf{g}}_{b}$, while the yaw component is set to 0 since we are not incorporating a compass/magnetometer. In addition, we derive the initial accelerometer bias as $\mathbf{b}_{a}=\mathbf{C}_{w}^{b}\left(\mathbf{C}_{b}^{w}\tilde{\mathbf{g}}_{b}-\mathbf{g}\right)$, thus taking into account the difference in magnitude between the expected and observed gravity vectors.
We initialize the covariance matrix mostly as 0, except for the subblocks related to $\mathbf{t_{r}^{b}}$, $\mathbf{b}_{a}$, $\mathbf{b}_{\omega}$, $\delta\boldsymbol{\theta}_{b}^{w}$ and $\delta\boldsymbol{\theta}_{r}^{b}$, whose diagonals we fill in according to supplied initial values.
V-B EKF propagation
We use a simplified inertial strapdown formulation that assumes movement to be linear instead of being perturbed by the current angular velocity reading of the gyroscope, due to the short time interval and small change in attitude (in effect equivalent to first order integration):
$\mathbf{p}$ $\leftarrow\mathbf{p}+\mathbf{v}t+\frac{1}{2}(\mathbf{C}^{w}_{b}(\mathbf{a}-\mathbf{b}_{a})+\mathbf{g})t^{2},$ (4)
$\mathbf{v}$ $\leftarrow\mathbf{v}+(\mathbf{C}^{w}_{b}(\mathbf{a}-\mathbf{b}_{a})+\mathbf{g})t,$ (5)
$\mathbf{q}$ $\leftarrow\mathbf{q}\otimes\exp\left(\frac{1}{2}(\boldsymbol{\omega}-\mathbf{b}_{\omega})t\right).$ (6)
The strapdown formulation is linearized in order to propagate the covariance matrix:
$\mathbf{P}\leftarrow\mathbf{F}\mathbf{P}\mathbf{F}^{T}+\mathbf{N}\mathbf{Q}\mathbf{N}^{T},$ (7)
where $\mathbf{F}$ is the linearized state transition matrix, $\mathbf{Q}$ is the noise covariance matrix, and $\mathbf{N}$ is the noise state transition matrix. These three matrices are sparsely populated: $\mathbf{F}=\mathbf{I}$ and $\mathbf{Q}=\mathbf{N}=\textbf{0}$, except for the following subblocks:
$\mathbf{F}_{p\,v}=\mathbf{I}t;\;\mathbf{F}_{p\,b_{a}}=-\frac{1}{2}\mathbf{C}^{w}_{b}t^{2};\;\mathbf{F}_{p\,\delta\theta_{b}^{w}}=-\frac{1}{2}[\mathbf{C}^{w}_{b}\mathbf{a}]_{\times}t^{2};$ (8)
$\mathbf{F}_{v\,b_{a}}=\mathbf{F}_{\delta\theta_{b}^{w}\,b_{\omega}}=-\mathbf{C}^{w}_{b}t;\;\mathbf{F}_{v\,\delta\theta_{b}^{w}}=-[\mathbf{C}^{w}_{b}\mathbf{a}]_{\times}t;$ (9)
$\mathbf{Q}_{v\,v}=\mathbf{I}\sigma^{2}_{v};\;\mathbf{Q}_{\theta\,\theta}=\mathbf{I}\sigma^{2}_{\theta};\mathbf{Q}_{a\,a}=\mathbf{I}\frac{\sigma^{2}_{a}}{t};\;$ (10)
$\mathbf{Q}_{\omega\,\omega}=\mathbf{I}\frac{\sigma^{2}_{\omega}}{t};\mathbf{Q}_{b_{a}\,b_{a}}=\mathbf{I}\sigma^{2}_{b_{a}}t;\;\mathbf{Q}_{b_{\omega}\,b_{\omega}}=\mathbf{I}\sigma^{2}_{b_{\omega}}t;$ (11)
$\mathbf{N}_{p\,a}=\frac{1}{2}\mathbf{C}^{w}_{b}t^{2};\;\mathbf{N}_{v\,a}=\mathbf{N}_{\delta\theta_{b}^{w}\,\omega}=\mathbf{C}^{w}_{b}t;$ (12)
$\mathbf{N}_{v\,v}=\mathbf{N}_{\delta\theta_{b}^{w}\,\theta}=\mathbf{N}_{b_{a}\,b_{a}}=\mathbf{N}_{b_{\omega}\,b_{\omega}}=\mathbf{I},$ (13)
where $\sigma_{v}$, $\sigma_{\theta}$, $\sigma_{a}$, $\sigma_{\omega}$, $\sigma_{b_{a}}$ and $\sigma_{b_{\omega}}$ are the standard deviations of the respective components of $\mathbf{w}$.
V-C EKF update and error state reset
Given observation $\mathbf{y}$, residual observation vector $\mathbf{r}$, observation covariance matrix $\mathbf{R}$ and linearized observation matrix $\mathbf{H}$ (so that $\mathbf{r}\approx\mathbf{y}-\mathbf{H}\mathbf{x}$), we can proceed to apply the Kalman filter equations (in Joseph form):
$\mathbf{K}=\mathbf{P}\mathbf{H}^{T}(\mathbf{H}\mathbf{P}\mathbf{H}^{T}+\mathbf{R})^{-1};\;\mathbf{L}=\mathbf{I}-\mathbf{K}\mathbf{H},$ (14)
$\mathbf{x}\leftarrow\mathbf{x}+\mathbf{K}\mathbf{r};\;\mathbf{P}\leftarrow\mathbf{L}\mathbf{P}\mathbf{L}^{T}+\mathbf{K}\mathbf{R}\mathbf{K}^{T}.$ (15)
As previously mentioned, rotations are expressed as error state in the EKF, and after an EKF update the nominal value of each $\delta\boldsymbol{\theta}$ becomes non-zero, making it necessary to incorporate the estimated error back into the state as follows:
$\mathbf{q}\leftarrow\delta\mathbf{q}\otimes\mathbf{q};\;\delta\mathbf{q}=\exp\left(\frac{1}{2}\delta\boldsymbol{\theta}\right),$ (16)
$\delta\boldsymbol{\theta}\leftarrow 0;\;\mathbf{P}\leftarrow\mathbf{G}\mathbf{P}\mathbf{G}^{T};\;\mathbf{G}_{\delta\theta\,\delta\theta}=\mathbf{R}\{\delta\mathbf{q}\},$ (17)
where $\mathbf{G}$ is the error reset matrix, and it is a sparse matrix equal to $\mathbf{I}$ except for every $\delta\theta\,\delta\theta$ subblock.
Finally, in order to prevent especially noisy observations from causing instability in the filter, we follow [29] and perform Mahalanobis outlier rejection, using a threshold calculated from a given percentile of the $\chi^{2}$ distribution.
V-D Radar egovelocity observation
Following previous works, we directly observe the robot’s velocity by leveraging the Doppler effect in measurements returned by the radar. We use the RANSAC-LSQ algorithm [29] to both filter out outlier (dynamic) points, and estimate the current radar-frame velocity of the robot, including a covariance matrix. The filtered point cloud is then fed to the Gaussian modeling and scan matching algorithm in the following subsection. The observation model used to perform the EKF update is as follows:
$\mathbf{r}=\mathbf{y}-\mathbf{C}_{b}^{r}\left([\boldsymbol{\omega}]_{\times}\mathbf{t}_{r}^{b}+\mathbf{C}^{b}_{w}\mathbf{v}\right);$ (18)
$\mathbf{H}_{:\,v}=\mathbf{C}_{b}^{r}\mathbf{C}_{w}^{b};\;\mathbf{H}_{:\,t_{r}^{b}}=\mathbf{C}_{b}^{r}[\boldsymbol{\omega}]_{\times};\;\mathbf{H}_{:\,b_{\omega}}=\mathbf{C}_{b}^{r}[\mathbf{t}_{r}^{b}]_{\times};$ (19)
$\mathbf{H}_{:\,\delta\theta_{b}^{w}}=\mathbf{C}_{b}^{r}\mathbf{C}_{w}^{b}[\mathbf{v}]_{\times};\;\mathbf{H}_{:\,\delta\theta_{r}^{b}}=\mathbf{C}_{b}^{r}[[\boldsymbol{\omega}]_{\times}\mathbf{t}_{r}^{b}+\mathbf{C}^{b}_{w}\mathbf{v}]_{\times}.$ (20)
V-E Scan matching observation
The Gaussian modeling and scan matching algorithms allow us to observe robot poses relative to previous keyframes. In other words, given a keyframe pose $\xi^{k}$ we can observe the change in translation and rotation that has occurred since said keyframe. We randomly initialize the pose particle swarm with a distribution $\mathbf{P}\thicksim\mathcal{N}(\xi^{x},\boldsymbol{\Sigma})$ where $\xi^{x}$ is the relative robot pose formed from the current nominal state of the EKF: $\xi^{x}=\{\mathbf{p},\mathbf{q}\}\ominus\xi^{k}$, and $\boldsymbol{\Sigma}$ is a covariance matrix indicating the desired dispersion of the particles for each of the 6 degrees of freedom (X/Y/Z/roll/pitch/yaw). Afterwards, we apply the scan matching algorithm and obtain a single optimized observation pose $\xi^{y}$, from which a residual pose can be obtained: $\xi^{r}=\xi^{y}\ominus\xi^{x}=\{\delta\mathbf{p},\delta\mathbf{q}\}$. Since the attitude error quaternion $\delta\mathbf{q}$ can be approximated as $\delta\mathbf{q}=\exp\left(\frac{1}{2}\delta\boldsymbol{\theta}\right)\approx[1\,\frac{1}{2}\delta\boldsymbol{\theta}^{T}]^{T}$ for very small errors, the attitude error vector can be approximated as $\delta\boldsymbol{\theta}\approx\frac{2}{\delta q_{w}}\delta\mathbf{q}_{v}$. This allows us to construct the residual vector $\mathbf{r}=[\delta\mathbf{p}^{T}\,\delta\boldsymbol{\theta}^{T}]^{T}$ to be fed to the EKF update equation – note that this vector contains X/Y/Z/roll/pitch/yaw components (in this order). The non-zero subblocks of the observation matrix $\mathbf{H}$ can be defined as such: $\mathbf{H}_{\delta p\,p}=\mathbf{H}_{\delta\theta\,\delta\theta}=\mathbf{C}_{w}^{k}$, where $k$ corresponds to the keyframe’s reference frame. We currently use a fixed observation covariance matrix $\mathbf{R}$, although we believe it is possible to leverage the optimized particle swarm returned by the scan matching algorithm to estimate this matrix.

### v5

@@ Abstract
4D millimeter-wave (mmWave) radars are sensors that provide robustness against adverse weather conditions (rain, snow, fog, etc.), and as such they are increasingly used for odometry and SLAM (Simultaneous Location and Mapping). However, the noisy and sparse nature of the returned scan data proves to be a challenging obstacle for existing registration algorithms, especially those originally intended for more accurate sensors such as LiDAR. Following the success of 3D Gaussian Splatting for vision, in this paper we propose a summarized representation for radar scenes based on global simultaneous optimization of 3D Gaussians as opposed to voxel-based approaches, and leveraging its inherent Probability Density Function (PDF) for registration. Moreover, we propose optimizing multiple registration hypotheses for better protection against local optima of the PDF. We evaluate our modeling and registration system against state of the art techniques, finding that our system provides richer models and more accurate registration results. Finally, we evaluate the effectiveness of our system in a real Radar-Inertial Odometry task. Experiments using publicly available 4D radar datasets show that our Gaussian approach is comparable to existing registration algorithms, outperforming them in several sequences.
@@ I Introduction
We introduce a new method for modeling and registering 4D radar point clouds using 3D Gaussians. Following the success of 3D Gaussian Splatting (3DGS) for vision [6], we propose the use of a collection of 3D Gaussians for modeling the geometry of a scene. As in [6] (and in contrast to NDT [1]), we simultaneously optimize all parameters of the model, including the position, scaling, rotation and correspondences between points and Gaussians. This approach is entirely featureless, and does not require any priors other than an initial set of centroids. Moreover, we propose a registration step that allows for multiple simultaneous hypotheses in order to add further robustness against local optimization minima. We evaluate our Gaussian approach against several established registration algorithms with a test suite independently evaluating the modeling and scan matching, as well as combined within a full Radar-Inertial Odometry (RIO) pipeline using an Extended Kalman Filter (EKF) on multiple datasets and radar types.
This paper is organized as follows: Section II reviews existing related literature, Sections III and IV describe the proposed Gaussian modeling and scan matching components respectively, Section V shows experimental results, and Section VI concludes the paper proposing future work.
@@ III-B Initialization
First, it is necessary to determine $N$, that is, the number of 3D Gaussians in the model. We propose choosing $N$ proportionally based on the number of points, so that richer scans can be represented with more Gaussians. The proportion should also consider the sparsity of the point clouds generated by the radar used. The model $\mathcal{G}$ is then initialized with all $\mathbf{s}_{j}$ set to 0 (indicating unit scale), and all $\mathbf{q}_{j}$ set to the unit quaternion. As for $\boldsymbol{\mu}_{j}$, we perform Bisecting K-Means clustering to find initial estimates for the centers of the Gaussians. Compared to voxel-based clustering used by NDT or VGICP [1, 15], this approach produces fewer spurious Gaussians from outlier points, as K-Means is biased towards higher density areas.
@@ III-C Optimization
IV Gaussian scan matching
Given a Gaussian model $\mathcal{G}$ and a point cloud $\mathcal{P}$, we define a registration process that estimates the pose of the robot $\xi$ with respect to $\mathcal{G}$. Moreover, we extend the usual approach to support the simultaneous optimization of multiple hypotheses centered around the current predicted position of the robot, in order to increase the robustness against local optima. At the end of the optimization, the hypotheses are judged against a scoring function, and the best hypothesis is selected as the output of the scan matching algorithm while the other hypotheses are discarded. Note that this is unlike a classical particle filter, which does persist the hypotheses throughout time.
@@ IV-A Definitions
We define the robot pose $\xi$ to be a member of $SE(3)$; that is, it includes a translational component $\mathbf{t}\in\mathbb{R}^{3}$ and rotational component $\mathbf{q}\in SO(3)$, with respect to some frame of reference. We consider that the uncertainty about the true robot pose follows some probability distribution: $\xi\thicksim\mathrm{P}$. Given a suitable model of $\mathrm{P}$, we can sample a swarm of $K$ pose particles $\hat{\xi}^{k}$, each representing a different hypothesis, and use an optimization algorithm that evaluates each particle against $\mathcal{G}$ to further refine these estimates until one of the particles reaches a suitable optimum.
@@ IV-B Optimization
  • We perform the usual Gauss-Newton update step for each particle $\hat{\xi}^{k}$, using the sums of squared Mahalanobis distances as the function to optimize, as in GICP. Each point contributes to the gradient according to the following weight:
    $w^{k}_{i}=\min\left(1,\frac{d_{\textrm{max}}}{d^{k}_{i}}\right);$ (2)
    where $d_{\textrm{max}}$ is a given Mahalanobis distance threshold.
@@
$\mathcal{L}_{k}=\frac{1}{M}\sum_{i}^{M}\min\left(d^{k}_{i},d_{\textrm{max}}\right),$ (3)
@@
V Experimental results
We use existing public 4D radar datasets for evaluation. In particular, we selected the NTU4DRadLM [8] and Snail-Radar [28] datasets, because together they cover multiple radar types (Oculii Eagle and ARS548), multiple robot platform types (handcart and car), and adverse conditions. In particular, NTU4DRadLM includes scans captured by an Oculii Eagle radar, and a VectorNav VN-100 IMU; while Snail-Radar uses both Oculii Eagle and ARS548 radars together with an Xsens MTi-3-DK IMU, and includes sequences recorded in harsher environment conditions.
We compare our Gaussian approach to three common established scan matching algorithms (NDT, GICP and VGICP). NDT uses the implementation provided by the well known PCL (Point Cloud Library), while GICP and VGICP use small_gicp [29] as the underlying implementation. In the case of Oculii Eagle scans, we set a target of 16 points per Gaussian; while for ARS548 scans, due to the much lower number of returned points, we set a target of 8 points per Gaussian. Moreover, these sparser scans trigger a known issue in PCL’s implementation of NDT that cause a program termination, so unfortunately NDT is excluded from experiments involving ARS548. Regarding our scan matching method, we provide results with 1, 8 and 128 particles wherever applicable. Other parameters include $d_{\mathrm{max}}$ = 4 for scan matching, and 5m/5° for particle dispersion. These parameters were chosen to suit the typical depth range/sparsity relation of a radar.
V-A Modeling experiments
In order to justify our modeling approach based on globally optimized 3D Gaussians versus existing voxel-based approaches, we compare in Fig. 1 the output of our method against NDT. To do this, we prepare a radar-based map of NTU4DRadLM’s cp sequence using the supplied ground truth, and model a section of said map with both NDT and our method. The positions and sizes of the Gaussians generated by NDT are clearly constrained and influenced by the voxel grid, while our approach generates better fitted Gaussians. Moreover, our method is able to preserve geometric details such as corners, and allocates a higher budget of Gaussians to areas with higher point density, resulting in overall better representation of the geometry.
Additionally, we include a video supplement. In the video we show several Gaussian models of the same map created using different points-per-Gaussian ratios. It can be seen that, regardless of said hyperparameter, our modeling method is able to generate a representation reasonably fitting the map. Different number of Gaussians are allocated to areas of the map depending on the density and complexity of the area, as well as the available budget. Moreover, when the ratio is higher (meaning fewer Gaussians), sparse regions with few (outlier) radar points do not receive any Gaussians, which we attribute to our use of K-Means clustering for initialization. Finally, we also include an example of the radar-inertial odometry pipeline on the same NTU4DRadLM cp sequence, compared to the ground truth trajectory.
Figure 2: Results of the sensibility study, showing failure rate (%), translation error (m) and rotation error (°). Lower is better. Green dots represent mean errors, with the numerical value shown below.